# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ Abstract
In this paper, we introduce a mesh-free two-level hybrid Tucker tensor format for approximation of multivariate functions, which combines the product Chebyshev interpolation with the ALS-based Tucker decomposition of the tensor of Chebyshev coefficients. It allows to avoid the expenses of the rank-structured approximation of function-related tensors defined on large spacial grids, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients. This leads to nearly optimal Tucker rank parameters which are close to the results for well established Tucker-ALS algorithm applied to the large grid-based tensors. These rank parameters inherited from the Tucker-ALS decomposition of the coefficient tensor can be much less than the polynomial degrees of the initial Chebyshev interpolant via function independent basis set. Furthermore, the tensor product Chebyshev polynomials discretized on a tensor grid leads to a low-rank two-level orthogonal algebraic Tucker tensor that approximates the initial function with controllable accuracy. It is shown that our techniques could be gainfully applied to the long-range part of the electrostatic potential of multi-particle systems approximated in the range-separated tensor format. Error and complexity estimates of the proposed methods are presented. We demonstrate the efficiency of the suggested method numerically on examples of the long-range components of multi-particle interaction potentials generated by 3D Newton kernel for large bio-molecule systems and lattice-type compounds.
Key words: Tensor numerical methods, Chebyshev interpolation, multilinear algebra, multi-particle electrostatic potentials, Tucker and canonical tensor formats.
AMS Subject Classification: 65F30, 65F50, 65N35, 65F10
@@ 1 Introduction
Tensor numerical methods (TNM) intrinsically reduce the numerical solution of multidimensional problems to one-dimensional calculations. In recent years, TNM are proven to be a powerful tool for solving numerically intensive multi-dimensional problems in scientific computing arising in computational quantum chemistry [33, 26], multi-dimensional dynamical problems, data science, stochastic simulations, optimal control problems, etc., see, e.g., [30]. The benefits of the TNM for $d$-dimensional problems are due to the reduction of both computational and storage complexity to merely linear scaling in $d$. Other approaches for optimizing numerical algorithms can be based on the concept of model reduction, see, e.g., [3, 6] and references therein.
@@
Tensor numerical methods provide the bridging of approximation theory for multi-dimensional functions and operators from one side and the modern multilinear algebra techniques from the other side. The construction of efficient TNM is guided by the trade-off between approximation accuracy and computational complexity of the chosen nonlinear rank-structured approximation algorithms, while taking into account the compatibility of the resultant tensor representations with available multilinear algebra. The total numerical complexity of tensor approximations is a sum of the high computational cost for (a) the algebraic operations of nonlinear low-rank approximation of multidimensional functions and operators in the traditional canonical (CP) [23], Tucker [49], tensor train (TT) [39], quantized tensor train (QTT) [29] or range-separated (RS) [4] tensor formats, and (b) the cost of function evaluations required for implementation of the tensor algorithms. Thereafter, the overall approximation accuracy can be controlled by tensor rank parameters as well as the precision of the chosen functional approximations (given the storage budget, accuracy is limited by the data regularity). A number of recent developments confirm efficiency of TNM in different research fields [47, 41, 35, 34, 44, 1, 43, 50].
@@
TNM are usually applied to grid-based rank-structured tensor approximation of target functions and operators, discretized on large $n^{\otimes d}$ spacial grids in $\mathbb{R}^{d}$ with $d=2,3,...$, for the numerical solution of large scale scientific computing problems, see [30, 26, 5]. Alternatively, classical mesh-free approximation methods for functions are based on their representation in problem-independent sets of functions, such as trigonometric functions (Fourier basis), (piecewise) polynomials, Sinc functions or Gaussians, [45, 48, 10, 38, 35, 18], as well as fast multipole methods [19].
This paper is motivated by recent advances in low-rank tensor techniques for tensor product Chebyshev interpolation [2, 16, 17, 22, 46]. Chebyshev polynomials [12] are among the most commonly used set of interpolating functions due to many beneficial features [8, 7, 38]. Error analysis for tensor product interpolation by Chebyshev polynomials was presented for example in [20]. MATLAB implementations of the Chebyshev interpolation method in 1D and 3D (chebfun and chebfun3) are available [48, 17, 22].
@@
The basic version of chebfun3 applies to the input functions presented in the form of canonical tensor format, thus using chebfun for 1D polynomial interpolation. Approximation of the target function in a separable form in 3D [22] is performed via a variant of the heuristic method adaptive cross approximation (ACA) [2] that combines ACA for the 2D slices with agglomeration of slices along the third spacial direction. Hence, the number of functional calls in such an approach is a multiple of polynomial degrees $m$, the CP rank of ACA approximation on each slice, $R_{cross}$, and the number of slices, $n_{slice}$. This procedure may reduce the number of functional calls under the condition $R_{cross}n_{slice}<m^{2}$. In turn, the direct tensor-product Chebyshev interpolant requires $m^{3}$ functional calls. The modification of Chebfun3 algorithm that reduces the number of functional calls was presented in [16].
In general, for functions of continuous arguments, one can apply mesh-free interpolation-type approximations in the form of a weighted sum of a given problem independent set of functions (not necessary orthogonal), where the corresponding weights require the values of the target function only at a few predefined interpolation points in a computational domain. This usually leads to asymptotically quasi-optimal accuracy for each particular input function having good smoothness characteristics. However, for functions with low regularity, satisfactory accuracy may require a rather large set of $m$ interpolating functions, say, with $m$ in the range of many hundreds if not thousands, which makes further use of such interpolants in scientific computing non-tractable. That is the main bottleneck for interpolation of multivariate functions by Chebyshev polynomials.
@@
In this paper, we introduce and analyze a novel method for approximating multivariate functions by using Chebyshev polynomial interpolation in $\mathbb{R}^{d}$ with $d=3$, which combines the tensor product Chebyshev polynomial interpolation in the global computational domain with the ALS-based Tucker decomposition of the rather small tensor of Chebyshev coefficients [^1] Using the non-orthogonal interpolating functions leads to small modification of the computational scheme.. This leads to the construction of mesh-free two-level Tucker tensor format for approximation of multivariate functions, either given explicitly or only available on a regular grid. It allows to avoid the expenses of the rank-structured approximation of function-related tensors defined on large spacial grids, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients, which leads to nearly optimal rank parameters as for the well established grid-based methods. The ranks inherited from the Tucker-ALS decomposition are almost optimal and can be much less than the polynomial degree of the initial Chebyshev interpolant which is chosen independent of the grid sizes in grid-based methods. Thus, we compute the nearly optimal Tucker decomposition of the 3D function without discretizing the function on the full grid in the computational domain, but only using its values at Chebyshev nodes. Finally, we can represent the function in Tucker format with the optimal $\varepsilon$-rank on an arbitrarily large 3D tensor grid in the whole computational domain by discretizing the Chebyshev polynomials on that grid. Note that in the case of functions with multiple cusps (or even stronger singularities), our method applies to the long-range component of the target function approximated in the range-separated (RS) tensor format [4].
@@
We underline that, given accuracy $\varepsilon>0$, our main goal in this paper is to construct numerical methods that allow to compute the mesh-free Tucker tensor approximation of the target function with almost optimal tensor rank ${\bf r}$ which is usually much smaller than the Chebyshev polynomials degree $m$ that guarantees the accuracy $\varepsilon$. In turn, the parameter $m$ should be much smaller than the grid size that is required for $\varepsilon$-accurate grid-based discretizations with the mesh-size $h$,
$|{\bf r}_{0}|\ll m\ll O(h^{-1}).$ (1.1)
@@
The Tucker rank of the output tensor is the most important parameter for the complexity characteristics of subsequent multilinear tensor calculus in respective applications.
@@
We present numerical experiments demonstrating the efficiency of the proposed techniques on nontrivial examples, including the computation of multi-particle electrostatic potentials arising in bio-molecular modelling and lattice type structure calculations in material science. In particular, it is gainfully applied to the long-range part of the many-particle electrostatic potentials in $\mathbb{R}^{3}$, obtained by the RS tensor format [4]. Numerical results indicate that the low-rank tensor representations can be computed with controllable accuracy such that the numerical cost depends weakly on the size of the many-particle system [4]. On the other hand, our numerics reproduce the earlier results obtained by the grid-based tensor numerical methods [31, 25]. In summary, we state the main benefits of the proposed method as follows:
@@
  • Fast and accurate quasi-optimal mesh-free Tucker tensor decomposition of regular functions with rather small number of Chebyshev polynomials.
  • Efficient quasi-optimal mesh-free Tucker decomposition of the long-range part of the multi-particle electrostatic potential, i.e., the large weighted sum of Newton kernels.
@@
  • In case of input function presented in the canonical (CP) tensor format our method needs only a small number of 1D-Chebyshev interpolation of canonical skeleton vectors.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the definition and construction of our mesh-free hybrid Chebychev-Tucker tensor representation. We discuss here transforming of the hybrid representation to the standard low-rank Tucker format. Section 3 presents the numerical schemes for the hybrid Chebyshev-Tucker approximation of trivariate functions, either given explicitly or only available on a regular grid. The corresponding theorems on error and complexity bounds of the Chebyshev-Tucker approximation conclude this section. Section 4 first outlines the techniques for grid-based low-rank tensor representations of multi-particle interaction potentials. We then discuss in details the Chebyshev-Tucker approximation of trivariate functions with singularities (the Newton kernel). We show efficiency of our methods for approximating the long-range part of the multi-particle electrostatic potentials of bio-molecular systems of different size, and present various numerical tests illustrating the performance of the proposed approach. Numerical results for application of the new method to lattice-type compounds are also presented in this section. Finally, Section 5 concludes the paper and outlines the future research directions.
@@ 2.1 Definition of hybrid tensor format
Recall that a tensor of order $d$ is defined as a real multidimensional array over a $d$-tuple index set
${\bf A}=[a_{i_{1},\cdots,i_{d}}]\equiv[a(i_{1},\cdots,i_{d})]\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}},$ (2.1)
with multi-index notation ${\bf i}=(i_{1},\cdots,i_{d})$, $i_{\ell}\in I_{\ell}:=\{1,\dots,n_{\ell}\}$. It is considered as an element of the Euclidean vector space $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$. To get rid of the exponential scaling in storage size and the consequent drawbacks, one can apply the rank-structured formatted separable approximations of multidimensional tensors.
The $R$-term canonical (CP) tensor format is defined by a finite sum of rank-1 tensors
${\bf U}_{R}=\sum_{k=1}^{R}\xi_{k}{\bf u}_{k}^{(1)}\otimes\cdots\otimes{\bf u}_{k}^{(d)},\quad\xi_{k}\in\mathbb{R},$ (2.2)
where ${\bf u}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are normalized vectors, and $R\in\mathbb{Z}_{+}$ is the canonical rank. We also call the vectors ${\bf u}_{k}^{(\ell)}$ the canonical vectors.
Recall that the rank-${\bf r}$, ${\bf r}=(r_{1},r_{2},\cdots,r_{d})$, orthogonal Tucker format is specified as the set of tensors ${\bf V}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ parametrized as follows,
${\bf V}=\sum_{\nu_{1}=1}^{r_{1}}\cdots\sum_{\nu_{d}=1}^{r_{d}}\beta_{\nu_{1},\cdots,\nu_{d}}{\bf v}_{\nu_{1}}^{(1)}\otimes\cdots\otimes{\bf v}_{\nu_{d}}^{(d)}\equiv\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\cdots\times_{d}V^{(d)},$ (2.3)
where $\{{\bf v}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}}$ is the set of orthonormal vectors for $\ell=1,\cdots,d$. Here $\times_{\ell}$ denotes the contraction product along the mode $\ell$ with the orthogonal matrices $V^{(\ell)}=[{\bf v}_{1}^{(\ell)},\cdots,{\bf v}_{r_{\ell}}^{(\ell)}]\in\mathbb{R}^{n_{\ell}\times r_{\ell}}$. The Tucker core tensor is denoted by $\bm{\beta}=[\beta_{\nu_{1},\cdots,\nu_{d}}]\in\mathbb{R}^{r_{1}\times\cdots\times r_{d}}$.
Next, we define the functional analog of the Tucker format, which we call the functional Tucker format.
@@
We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $[-1,1]^{3}$ is represented in the functional rank-${\bf r}$ Tucker format if it allows the factorization
$f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}F^{(1)}(x_{1})\times_{2}F^{(2)}(x_{2})\times_{3}F^{(3)}(x_{3}),\quad F^{(\ell)}:[-1,1]\mapsto\mathbb{R}^{r_{\ell}},$ (2.4)
where $\bm{\beta}=\{\beta_{\nu_{1},\nu_{2},\nu_{3}}\}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is the core tensor and $F^{(\ell)}(x_{\ell})$ is the Tucker factor function defined as a vector-valued function with functional entries, i.e., $F^{(\ell)}(x_{\ell})=(f_{1}(x_{\ell}),\cdots,f_{r_{\ell}}(x_{\ell}))$ for $f_{\nu_{\ell}}:[-1,1]\mapsto\mathbb{R}$, $\nu=1,\cdots,r_{\ell}$, $\ell=1,2,3$.
In the following definition we distinguish the discrete and functional two-level hybrid tensor formats formulated for the three-dimensional case, $d=3$.
@@
(A) (Functional case) If the core tensor $\bm{\beta}$ in Eq. 2.4 is in turn a rank-${\bf r}_{0}$ Tucker tensor,
$\bm{\beta}=\bm{\beta}^{0}\times_{1}G^{(1)}\times_{2}G^{(2)}\times_{3}G^{(3)},\quad G^{(\ell)}\in\mathbb{R}^{r_{\ell}\times r_{\ell}^{0}},$ (2.5)
then we call $f$ is represented in the two-level hybrid functional rank-$({{\bf r}_{0},\bf r})$ Tucker format.
(B) (Algebraic case). We say that tensor ${\bf A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is parametrized in the two-level hybrid rank-$({{\bf r}_{0},\bf r})$ Tucker-Tucker algebraic format if it admits the representation
${\bf A}=\bm{\beta}\times_{1}A^{(1)}\times_{2}A^{(2)}\times_{3}A^{(3)},\quad\mbox{with}\quad A^{(\ell)}=\{{\bf a}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}\times r_{\ell}},\quad\ell=1,2,3,$ (2.6)
@@
Furthermore, if the functional entries of $F^{(\ell)}$ in Eq. 2.4 are specified by the Chebyshev polynomials, we say that the function $f(x_{1},x_{2},x_{3})$ is represented in the (two-level hybrid) Chebychev-Tucker format, shortly ChebTuck format. See Definition 2.3 below.
@@
Let $T_{k}(x)=\cos(k\arccos(x))$ denote the $k$-th Chebyshev polynomial of the first kind on $[-1,1]$, which forms a sequence of orthogonal polynomials on $[-1,1]$ with respect to the weight $w(x)=1/\sqrt{1-x^{2}}$. Let $T_{0:k-1}(x)=(T_{0}(x),\cdots,T_{k-1}(x))$ denote the vector-valued function of the first $k$ Chebyshev polynomials. We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $[-1,1]^{3}$ is represented in the ChebTuck-$(\bf{m},\bf{r})$ format if it allows the factorization
$f(x_{1},x_{2},x_{3})={\bf C}\times_{1}T_{0:m_{1}-1}(x_{1})\times_{2}T_{0:m_{2}-1}(x_{2})\times_{3}T_{0:m_{3}-1}(x_{3}),$ (2.7)
@@
${\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}},\quad V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}.$ (2.8)
For the practical application of the presented hybrid Tucker-type tensor representation in scientific computing, one requires the implementation of the standard multilinear algebra operations within arising tensor calculus. This means that tensors in hybrid format might be easily transformed to the standard Tucker-type format when required. This issue will be considered in the following Section 2.2.
2.2 Transforming hybrid representation to the standard low-rank Tucker format
The following Proposition 2.4 proves that any tensor parametrized as the two-level orthogonal Tucker format with ranks $({\bf r}_{0},{\bf r})$ can be converted to the standard Tucker tensor with the reduced (quasi-optimal) rank parameters $|{\bf r}_{0}|<|{\bf r}|$. This allows the traditional multilinear algebra operations with the tensors in the hybrid format. Moreover, the resultant Tucker tensor can be transformed to a canonical format with small rank.

### v2

@@ Abstract
We introduce and analyze a mesh-free two-level hybrid Tucker tensor format for approximating multivariate functions, which combines the product Chebyshev interpolation with the alternating least-squares (ALS) based Tucker decomposition of the tensor of Chebyshev coefficients. This construction allows to avoid the expensive rank-structured grid-based approximation of function-related tensors on large spatial grids, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients. Thus, we can compute the nearly optimal Tucker decomposition of the 3D function with controllable accuracy $\varepsilon>0$ without discretizing the function on the full grid in the domain, but only using its values at small set of Chebyshev nodes. Finally, we can represent the function in the algebraic Tucker format with optimal $\varepsilon$-rank on an arbitrarily large 3D tensor grid in the computational domain by discretizing the Chebyshev polynomials on that grid. The rank parameters of the Tucker-ALS decomposition of the coefficient tensor can be much smaller than the polynomial degrees of the initial Chebyshev interpolant obtained via a function independent polynomial basis set. It is shown that our techniques can be gainfully applied to the long-range part of the singular electrostatic potential of multi-particle systems approximated in the range-separated tensor format. We provide error and complexity estimates and demonstrate the computational efficiency of the proposed techniques on challenging examples, including the multi-particle electrostatic potential for large bio-molecular systems and lattice-type compounds.
keywords:
Tensor numerical methods, Tucker and canonical tensor formats, Chebyshev polynomial interpolation, multilinear algebra, multi-particle electrostatic potentials.
{MSCcodes}
65F30, 65F50, 65N35, 65F10
@@ 1 Introduction
Tensor numerical methods (TNM) intrinsically simplify the numerical solution of multi-dimensional problems to one-dimensional calculations. In recent years, TNM are proven to be a powerful tool for solving numerically intensive multi-dimensional problems in scientific computing arising in computational quantum chemistry [34], multi-dimensional steady state and dynamical PDE-driven problems, data science, stochastic simulations, optimal control problems, etc., see, e.g., [1, 37] for a detailed discussion and related references. The benefits of TNM for $d$-dimensional problems are due to the reduction of both computational and storage complexity to merely linear scaling in $d$ based on the representation of functions and operators in low-parametric rank structured tensor formats.
@@
TNM originated as bridging of approximation theory for multi-dimensional functions and operators from one side and the modern multilinear algebra techniques from the other side. The construction of efficient TNM is guided by the trade-off between approximation accuracy and computational complexity of the chosen rank-structured approximation algorithms, while taking into account the compatibility of the resultant tensor representations with available multilinear algebra. The total numerical complexity of tensor approximations is a sum of the high computational (and storage) cost for (a) the algebraic operations needed for the low-rank approximation of multi-dimensional functions and operators in the conventional canonical (CP) [30], Tucker [60], and rather recent tensor train (TT) [48], quantized TT (QTT) [36] or range-separated (RS) [6] tensor formats, and (b) the cost of function evaluations required for implementation of the tensor algorithms. Thereafter, the overall approximation accuracy can be controlled by rank parameters as well as the precision of the chosen functional approximations (given the storage budget, accuracy is limited by the data regularity). A number of recent developments confirm efficiency of TNM in different research fields [2, 42, 55, 44, 43, 52, 51, 61].
@@
TNM are usually applied to grid-based rank-structured tensor approximation of target functions and operators discretized on large $n^{\otimes d}$ spatial grids in $\mathbb{R}^{d}$, $d\geq 2$, for the numerical solution of large scale scientific computing problems [37, 34, 7]. In this direction the sparse grid methods, stochastic homogenization techniques and UQ approaches also provide powerful computational tools for special classes of problems. Alternatively, classical mesh-free approximation methods for functions are based on their representation in problem-independent basis sets, such as trigonometric functions (Fourier), (piecewise) polynomials, sinc functions or Gaussians [53, 59, 13, 47, 44, 24]. The fast multipole method provides another powerful example of a mesh-free approach for certain kernel-based problems [25]. The other well-developed computational technique is based on the general principle of model reduction [4, 8].
@@
This paper is partially motivated by recent advances in tensor techniques for 3D Chebyshev interpolation [3, 20, 22, 29, 54]. Chebyshev polynomials [16] are among the most commonly used set of interpolating functions due to many beneficial features [10, 9, 47]. Error analysis for tensor product interpolation by Chebyshev polynomials was presented for example in [27]. MATLAB implementations of the Chebyshev interpolation method up to 3D are available within the chebfun package [59, 22, 29, 57, 58, 56, 12]. The basic version of the trivariate chebfun3 [29] applies to ${\bf T}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ represented in Tucker format, which contains the values of the input function at Chebyshev nodes, thus using only chebfun for 1D polynomial interpolation. Approximation of the target tensor $\bf T$} in Tucker format is performed via a variant of the heuristic method adaptive cross approximation (ACA) [3] that combines ACA for the 2D slices with agglomeration of slices along the third spatial direction. Hence, the number of functional calls in this approach is $\mathcal{O}(mr^{2})$ and may be beneficial for low-rank smooth functions. Here $m$ is the maximum Chebyshev degrees in the three dimensions, and $r$ is the Tucker rank. In contrast, the direct tensor product Chebyshev interpolant requires $m^{3}$ function calls. A modification of chebfun3 that reduces the number of function calls was presented in [20].
While powerful, these interpolatory methods face a significant challenge. For functions with sufficient smoothness, this approach can achieve quasi-optimal accuracy with a relatively small basis set. However, for functions with low regularity or localized features, such as those with sharp gradients or singularities, achieving a satisfactory accuracy may require a very large number of interpolating functions (i.e., a high polynomial degree $m$), which makes further use of such interpolants in scientific computing non-tractable. That is the main bottleneck for approximation of multivariate functions by Chebyshev polynomials in large scale scientific computing involving low regularity functions.
@@
In this paper, we introduce and analyze a novel method that directly addresses this challenge, which combines the tensor product Chebyshev polynomial interpolation in the computational domain with the ALS-based Tucker decomposition of the rather small tensor of Chebyshev coefficients [^1] Using non-orthogonal interpolating functions leads to small modification of the computational scheme., termed Chebyshev-Tucker (ChebTuck) format. This leads to the construction of mesh-free two-level Tucker tensor format for approximation of multivariate functions, either given explicitly in analytic form or only available in the form of discrete grid-based representations. It allows to avoid the expenses of the rank-structured approximation of function-related tensors defined on fine spatial grids in $\mathbb{R}^{3}$, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients, which leads to nearly optimal rank parameters as for the well established grid-based methods. The $\varepsilon$-ranks inherited from the Tucker-ALS decomposition are almost optimal and can be much smaller than the polynomial degree, $m$, of the initial Chebyshev interpolant which, in turn, is chosen independent of the grid sizes in grid-based methods. Thus, we compute the nearly optimal Tucker decomposition of the 3D function without discretizing the function on the full grid in the computational domain, but only using its values at Chebyshev nodes. Note that in the case of functions with multiple cusps (or even stronger singularities), our method applies to the long-range component of the target function approximated in the range-separated (RS) tensor format [6].
@@
We underline that, given accuracy $\varepsilon>0$, our main goal in this paper is to construct numerical methods that allow to compute the mesh-free Tucker tensor approximation of the target function with almost optimal tensor rank ${\bf r}=(r_{1},r_{2},r_{3})$ which is usually much smaller than the Chebyshev polynomials degree $m$ that guarantees the accuracy $\varepsilon$ of the Chebyshev interpolant. In turn, the parameter $m$ is supposed to be much smaller than the grid size $n$ that is required for $\varepsilon$-accurate grid-based discretization with the mesh-size $h$,
(1) $|{\bf r}|\ll m\ll n=O(h^{-1}).$
@@
The resulting Tucker rank is the most important parameter governing the complexity of subsequent multilinear algebra operations.
@@
We present numerical experiments demonstrating the efficiency of the proposed techniques on nontrivial examples, including the computation of multi-particle electrostatic potentials arising in bio-molecular modeling and lattice type structures calculations in material science. In particular, it is gainfully applied to the long-range part of the many-particle electrostatic potentials in $\mathbb{R}^{3}$, obtained by the RS tensor format [6]. Numerical results indicate that the low-rank tensor representations can be computed with controllable accuracy, at a cost that depends only weakly on the number of particles [6]. In summary, we state the main benefits of the proposed method as follows:
@@
  • Fast and accurate mesh-free decomposition: The method yields a quasi-optimal mesh-free Tucker decomposition for both regular functions and the long-range part of functions with singularities, such as the multi-particle electrostatic potential, by using rather small number of Chebyshev polynomials.
@@
  • Efficiency for CP-formatted input: In case of input functions presented in the canonical (CP) tensor format, our method needs only a small number of 1D Chebyshev interpolations of skeleton vectors, followed by a reduced HOSVD [39, 35] of the Chebyshev coefficient tensor, thus resulting in a complexity that scales linear in the number of dimensions $d$.
  • Compatibility with grid-based methods: By discretizing the Chebyshev polynomials in the final ChebTuck format, one obtains a standard low-rank Tucker tensor on any desired fine grid. This avoids the direct, and often prohibitively expensive, decomposition of a large grid-based tensor in $\mathbb{R}^{n\times n\times n}$, while reproducing the results of established grid-based methods [39, 33].
@@
The remainder of the paper is organized as follows. In Section 2, we formally define the hybrid Chebychev-Tucker tensor representation and discuss its transformation to the standard Tucker format. Section 3 presents the numerical schemes for constructing the approximation of trivariate functions, either given explicitly or only available on a regular grid, and provides corresponding error and complexity bounds. Section 4 first outlines the well-developed techniques for grid-based low-rank tensor representations of multi-particle interaction potentials. We then discuss in detail the Chebyshev-Tucker approximation of trivariate functions with singularities (the Newton kernel). We show efficiency of our methods for approximating the long-range part of the multi-particle electrostatic potentials of bio-molecular systems and lattice-type compounds of different sizes, and present various numerical tests illustrating the asymptotic performance of the proposed approach. Finally, Section 5 concludes the paper and outlines potential future research directions.
@@ 2.1 Definition of hybrid tensor format
Recall that a tensor of order $d$ is defined as a real multi-dimensional array over a $d$-tuple index set ${\bf A}=[a_{i_{1},\cdots,i_{d}}]\equiv[a(i_{1},\cdots,i_{d})]\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}},$ with multi-index notation ${\bf i}=(i_{1},\cdots,i_{d})$, $i_{\ell}=1,\cdots,n_{\ell}$. To get rid of the exponential scaling in storage size, one can apply the rank-structured formatted separable approximations of tensors. For example, the $R$-term canonical (CP) tensor format is defined by a finite sum of rank-1 tensors
(2) ${\bf U}_{R}={\sum}_{k=1}^{R}\xi_{k}{\bf u}_{k}^{(1)}\otimes\cdots\otimes{\bf u}_{k}^{(d)},\quad\xi_{k}\in\mathbb{R}_{+},$
where $R\in\mathbb{N}_{+}$ is the canonical rank, ${\bf u}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are normalized vectors, also known as the canonical (skeleton) vectors. The rank-$\bf r$}, ${\bf r}=(r_{1},\cdots,r_{d})$, orthogonal Tucker format is specified as the set of tensors ${\bf V}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ parametrized as
(3) ${\bf V}=\sum_{\nu_{1}=1}^{r_{1}}\cdots\sum_{\nu_{d}=1}^{r_{d}}\beta_{\nu_{1},\cdots,\nu_{d}}{\bf v}_{\nu_{1}}^{(1)}\otimes\cdots\otimes{\bf v}_{\nu_{d}}^{(d)}\equiv\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\cdots\times_{d}V^{(d)},$
where $\{{\bf v}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}}$ is a set of orthonormal vectors for $\ell=1,\cdots,d$. Here, $\times_{\ell}$ denotes the contraction along the $\ell$-th mode of the Tucker core $\bm{\beta}\in\mathbb{R}^{r_{1}\times\cdots\times r_{d}}$ with the second mode of the orthogonal matrices $V^{(\ell)}=[{\bf v}_{1}^{(\ell)},\cdots,{\bf v}_{r_{\ell}}^{(\ell)}]\in\mathbb{R}^{n_{\ell}\times r_{\ell}}$. Here and after we consider $d=3$ to simplify the presentation since (1) the extension to higher dimensions is straightforward, and (2) Tucker is not the recommended format for high dimensions due to the exponential scaling of the storage of the core tensor.
Next, we define the functional analog of the Tucker format and distinguish the discrete and functional two-level hybrid tensor formats in Definitions 2.1 and 2.2.
@@
We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $1,1]^{3}$ is represented in the functional rank-$\bf r$} Tucker format if it allows the factorization
(4) $f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}F^{(1)}(x_{1})\times_{2}F^{(2)}(x_{2})\times_{3}F^{(3)}(x_{3}),\quad F^{(\ell)}:[-1,1]\mapsto\mathbb{R}^{r_{\ell}},$
where $\bm{\beta}=[\beta_{\nu_{1},\nu_{2},\nu_{3}}]\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is the core tensor and $F^{(\ell)}(x_{\ell})$ is the Tucker factor function defined as a vector-valued function with functional entries, i.e., $F^{(\ell)}(x_{\ell})=(f_{1}^{(\ell)}(x_{\ell}),\cdots,f_{r_{\ell}}^{(\ell)}(x_{\ell}))$ for $f_{\nu_{\ell}}^{(\ell)}:[-1,1]\mapsto\mathbb{R}$, $\nu_{\ell}=1,\cdots,r_{\ell}$, $\ell=1,2,3$.
@@
(A) (Functional case) We say that $f$ is in the two-level hybrid functional rank-$({{\bf r}_{0},\bf r})$ Tucker format if the core tensor $\bm{\beta}$ in Eq. 4 is in turn a rank-${\bf r}_{0}$ Tucker tensor,
(5) $\bm{\beta}=\bm{\eta}\times_{1}G^{(1)}\times_{2}G^{(2)}\times_{3}G^{(3)},\quad G^{(\ell)}\in\mathbb{R}^{r_{\ell}\times r_{\ell}^{0}}.$
(B) (Algebraic case). We say that tensor ${\bf A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is parametrized in the two-level hybrid rank-$({{\bf r}_{0},\bf r})$ Tucker-Tucker algebraic format if it admits
(6) ${\bf A}=\bm{\beta}\times_{1}A^{(1)}\times_{2}A^{(2)}\times_{3}A^{(3)},\quad\mbox{with}\quad A^{(\ell)}=\{{\bf a}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}\times r_{\ell}},\quad\ell=1,2,3,$
@@
Furthermore, let $T_{k}(x):=\cos((k-1)\arccos(x))$ denote the (k1)(k-1)-st Chebyshev polynomial of the first kind, which forms a sequence of orthogonal polynomials on [1,1][-1,1] with respect to the weight $w(x)=1/\sqrt{1-x^{2}}$. If the functional entries of $F^{(\ell)}$ in Eq. 4 are specified by the Chebyshev polynomials, we say that $f$ is represented in the (two-level hybrid) Chebyshev-Tucker format, shortly ChebTuck format.
@@
Let $T_{1:k}(x)=(T_{1}(x),\cdots,T_{k}(x))$ denote the vector-valued function of the first $k$ Chebyshev polynomials. We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $1,1]^{3}$ is represented in the ChebTuck-$(\bf{m},\bf{r})$ format if it allows the factorization
(7) $f(x_{1},x_{2},x_{3})={\bf C}\times_{1}T_{1:m_{1}}(x_{1})\times_{2}T_{1:m_{2}}(x_{2})\times_{3}T_{1:m_{3}}(x_{3}),$
@@
(8) ${\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}},\quad V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}.$
For the practical application in scientific computing the presented hybrid Tucker-type tensor representation might be easily transformed to the standard Tucker-type format when required.
2.2 Transforming hybrid representation to the standard Tucker format
The following Proposition 2.4 proves that any tensor parametrized in the two-level orthogonal Tucker format with ranks $({\bf r}_{0},{\bf r})$ can be converted to the standard Tucker tensor with an reduced (quasi-optimal) rank $|{\bf r}_{0}|<|{\bf r}|$. This allows the traditional multilinear algebra operations with tensors in the hybrid format.